OpenStudy (anonymous):

Solve the separable differential equation 7x-4ysqrt(x^2+1)dy/dx=0 Subject to the initial condition: y(0)=3

OpenStudy (anonymous):

\[3x-4y\sqrt{x^2+1}\frac{dy}{dx}=0 \] Like that?

OpenStudy (anonymous):

\[\frac{3x}{\sqrt{x^2+1}} dx=4ydy\] Where do you get stuck here?

OpenStudy (anonymous):

it took the previous question with a negative ivp... \[7x-4y \sqrt{x^2+1}\frac{ dy }{dx }=0 \] y(0)=3

OpenStudy (anonymous):

Pardon me, not sure why I turned the \(7x\) into a \(3x\) in case the question didn't update ^^, however yes, it's separable and easy to integrate. The initial condition will only help you find the yet unknown constant.

OpenStudy (anonymous):

i'm getting \[y(x)=\sqrt{\frac{ \sqrt{(3x \sqrt{x^2+1}+18)} }{ 2 }}\]

OpenStudy (anonymous):

it's saying the answer isn't correct, I don't know where I went wrong

OpenStudy (anonymous):

Lets look at the left hand side first, because it's the more complicated integral. \[\Large \int \frac{7x}{\sqrt{x^2+1}}dx \] this integral can easy be solved with an u substitution: u=x^2+1 \[\Large \frac{7}{2}\int \frac{1}{\sqrt{u}}du=7\sqrt{x^2+1} \]

OpenStudy (anonymous):

and for the right hand side we get 2y^2

OpenStudy (anonymous):

You can now put your initial condition on either side and solve for it.

OpenStudy (anonymous):

\[\Large 7\sqrt{x^2+1}=2y^2+C \]

OpenStudy (anonymous):

\[y(x)=\sqrt{\frac{ 7\sqrt{x^2+1}-7\sqrt{1}-18 }{ 2 }}\] is what i'm getting now

OpenStudy (anonymous):

\[\Large 7\sqrt{x^2+1}=2y(x)^2+C \] \[\Large y(0)=3 \] \[\Large 7\sqrt{0^2+1}=2\cdot3^2+C \] such that: \[C=-11 \]

OpenStudy (anonymous):

which should then give me \[y(x)=\sqrt{\frac{ 7\sqrt{x^2+1}+11 }{ 2 }}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for some reason it's not accepting that as the answer, I worked it out a few times and I keep getting the same answer now, i'm hoping the answer was saved wrong in the system

OpenStudy (anonymous):

It's right ;) well done

OpenStudy (anonymous):

thank you!